A proton of mass m_p = 1.67 \times 10^{-27}~kgm ​p ​​ =1.67×10 ​−27 ​​ kg and a charge of q_p = 1.60 \times 10^{-19} ~Cq ​p ​​ =1.60×10 ​−19 ​​ C is moving through vacuum at a constant velocity of 10,000 ~m/s10,000 m/s directly to the east when it enters a region of uniform electric field that points to the south with a magnitude of | \vec{E} | = ∣ ​E ​⃗ ​​ ∣=2.60e+3~N/C N/C . The region of uniform electric field is 5 mm wide in the east-west direction. How far (in meters) will the proton have been deflected towards the south by the time it exits the region of uniform electric field. You may neglect the effects of friction and gravity, and assume that the electric field is zero outside the specified region.
Answer #1

given

mp = 1.67 x 10-27 kg

p ​​ = 1.60 x 10−19 ​​ C

a constant velocity of 10,000 m/s

uniform electric field is 5 mm = 5 x 10-3 m

∣ ​E ​⃗ ∣ = 2.60 x 103 N/C

the time taken to reach between points is

10000 x t1 = 5 x 10-3

t1 = 5 x 10-7 sec

the acceleration of the proton is a = F / m

= q E / m

= 1.6 x 10-19 x 2600 / 1.67 x 10-27

a = 2.49 x 1011 m/sec2

now to find the distance

S = 0 x t1 + 1/2 x a t12

= 0 + 1/2 x 2.49 x 1011 x t12

= 1/2 x 2.49 x 1011 x (5 x 10-7)2  

S = 0.031125 m

far (in meters) will the proton have been deflected towards the south by the time it exits the

region of uniform electric field is S = 0.031125 m